# Can AR Embedded Visualizations Foster Appropriate Reliance on AI in Spatial Decision-Making? A Comparative Study of AR X-Ray vs. 2D Minimap

### v3

@@ 3.1. Task Abstraction and Definition
Spatial decision-making has long been studied under the broader umbrella of wayfinding in spatial cognition (Emo2012WayfindingAS; Wiener2009TaxonomyOH). Foundational work by Siegel and White (Siegel1975TheDO) conceptualizes wayfinding as a decision process involving three forms of spatial knowledge: landmark, route, and survey. Motivated by real-world applications, our study particularly focuses on target selection (landmark). In tasks such as exploring a building, security surveillance, and emergency response, spatial decision-making often involves identifying or selecting a target before proceeding with an action like navigation or intervention. Although AR systems have frequently been investigated for supporting the action phase such as route guidance (XUARteambasedsearch), their potential to enhance the preceding decision-making step remains underexplored.
@@
  • Space - Large-scale Indoor Environment: These tasks often take place in physical spaces larger than the immediate space around a person such as buildings or open areas (Giuliano2013IndoorLS; Guerrieri2006RFIDassistedIL). Tversky (Barbara2003StructureOfMentalSpace) characterize such spaces as space of navigation. In this study, we focus on indoor building environments, given their prevalence across numerous application domains.
  • Time - Sensitive: Many high-stakes spatial tasks impose stringent time constraints, requiring users to make optimal decision as soon as possible (nydegger2011post). In this study, we focus on time-pressure scenarios that particularly benefit from decision support systems.
  • Data - Static and Dynamic: Appropriate target selection relies on both static data like building level layouts, and dynamic updates from sensors like walking people, occupant counts, and hazard alerts to evaluate each target (Adib2013SeeThroughWallsWithWifi).
  • Decision - Multi-Criteria: These tasks frequently require decisions among multiple targets or locations distributed over rooms, floors, or areas (Launder2014ASI). Success often involves balancing or prioritizing several factors, such as response time, travel distance, and urgency (penney2022threat).
@@
X-Ray. Given our task focuses on large-scale indoor environments, we select AR X-ray (Figure 1 Left) as our embedded visualizations, which overlays a 1:1 scale 3D map on the real-world. It allows users see the interior layouts or objects behind walls, as if they had “X-ray” vision (Livingston2013PursuitO), which has been widely used in applications like navigation (XUImroveIndoorWayfinding) and construction (Soria2018AugmentedAV). For our study, we follow previous similar research (XUImroveIndoorWayfinding) and portray the building’s indoor layout, along with multiple targets (represented as cubes), and their associated data (the number of people in line) using virtual avatars.
Alternatives. World-in-Miniature (WiM) (Stoakley1995WIM; Danluk2021WiMs) provides a three-dimensional representation of the environment and can be implemented on flat-screen or AR displays. However, WiM typically relies heavily on interactions such as panning, tilting, zooming, or rotation to fully leverage their advantages. Because our current investigation prioritizes evaluating the impact of the visual representations themselves (without the influence of interactive complexity), we intentionally avoided such interactions in our study. Thus, we reserve exploration of interactive WiM and other advanced visualization methods for future research.
@@ 3.3. Visual Decision Supports
A recent survey by Oral et al. (Oral_2023_ACriticalInquiry) highlights a broad spectrum of visual decision support tools, ranging from basic information visualizations to advanced AI-based suggestions (Lawless2024IWantItThatWay). For this study, we specifically choose to implement two fundamental yet representative forms of decision support across both the Minimap and X-ray:
@@
  • AI-Suggestions that highlight a single optimal target based on predefined criteria, with no accompanying explanation. Similar to all AI systems, suggestions may not always be accurate. We detail our adjustment of the performance of the AI based on prior works (Swaroop2024AccuracyTime) and pilot studies in Sec. 4.7.
Following practices established in prior research (vasconcelos2023explanationsreduceoverrelianceai), we deliberately constrain our decision support to these fundamental features for several reasons: First, target and real-time data visualization represent core functionalities broadly used in existing decision-support tools. Second, simple AI suggestions enable direct investigation of user reliance on AI assistance under different visualizations, without the additional cognitive complexity and potential confounding effects introduced by detailed explanations (Jiang2018ToTrustOrNotToTrust).
Although more advanced support, such as comparative visualizations, AI-generated explanations, and uncertainty displays, may offer added benefits, we intentionally focus on these fundamental forms to isolate the role of embedded visualizations in spatial decision-making and AI reliance. This controlled setup provides a baseline for future studies exploring richer decision-support designs.
@@ 4.2. Spatial Arrangement
[…]
@@
[…]
@@
Since the values of $d$ were fixed by the spatial arrangement of the targets, we manipulated the queue lengths $n$ to differentiate target optimality. Following prior research (Swaroop2024AccuracyTime), each trial was designed to include intentionally varied options: one optimal, two suboptimal, and one worst choice, differing by at least 10 seconds in $T_{\text{coffee}}$. This setup allowed us to evaluate potential participant over-reliance on AI suggestions. Detailed calculations are provided in the supplementary materials.
@@ 4.3. Implementation and Apparatus
Following similar work (2021ToTrustOrToThink), a simulated AI was used to provide decision suggestions in applicable trials. Details of the simulated AI setup are provided in Sec. 4.7. We developed the minimap based on high-precision CAD models provided by the building manager. The AR visualizations was developed using ARKit  [^3] https://developer.apple.com/augmented-reality/arkit/ and AR Mesh Manager  [^4] https://docs.unity3d.com/Packages/com.unity.xr.arkit@6.2/manual/arkit-meshing.html to scan the building and construct a 3D digital twin in Unity 6000.0.29f1, with ARAnchor Manager  [^5] https://docs.unity3d.com/Packages/com.unity.xr.arfoundation@6.2/manual/features/anchors/aranchormanager.html ensuring proper registration of the digital model with the physical environment. The system is available at our public repository https://github.com/demoPlz/ARAIReliance. All trials were completed using an Apple Vision Pro to eliminate device-related differences and provided a high-fidelity immersive experience.
@@ 4.6. Measures
For every trial, we documented and calculated the following measures based on prior research (Swaroop2024AccuracyTime; Salimzadeh2024dealWithUncertainty):
  • Decision Accuracy: We defined decision accuracy as the average point score across trials in each condition, ranging from 0 to 1. Following previous research (Swaroop2024AccuracyTime), participants earned points according to their choice quality: optimal choice (1 point), suboptimal choices (0.5 points), and the worst choice (0 points). Each trial contained exactly one optimal target, two suboptimal targets, and one worst target.
  • AI Reliance: Following previous studies (yang2020visual), we categorized participant reliance on AI as either appropriate or inappropriate (Figure 6). We classified reliance as appropriate when participants either followed correct AI suggestions or overrode incorrect ones with better choices. Inappropriate reliance included: (1) over-reliance: accepting suboptimal or worst suggestions; and (2) under-reliance: rejecting correct suggestions in favor of a worse option.
@@
  • Point Error: Similar to prior research (hu2021comparative; warden2022visual), we assessed whether participants correctly pointed to the spatial location of their selected target based on first-person video recordings. Two co-authors independently coded each response as 0 (correct) or 1 (incorrect) to ensure inter-rater reliability. Inter-rater reliability was evaluated via Cohen’s Kappa ($\kappa=0.898$), suggesting almost perfect agreement according to Landis and Koch’s interpretation (landis1977measurement). Discrepancies were discussed and resolved through consensus. This measure indicates if participants actually knew the spatial location of their selected target.
@@ 4.7. Using Simulated AI
Similar to prior works (Swaroop2024AccuracyTime; 2021ToTrustOrToThink), we used a simulated AI with an average accuracy of 0.75 to ensure controlled experimentation. In each AI condition, the AI suggested the optimal target in five trials, a suboptimal target in two trials, and the worst target in one trial (Figure 5). This setup yields an average accuracy of 0.75 = $\frac{5\times 1+2\times 0.5+1\times 0}{8}$. The order of these suggestions was randomized and counterbalanced across participants, ensuring that optimal, suboptimal, and worst suggestions were evenly distributed across all trials and participants.
Mitigating Effects of Trust in AI. Following previous work (2021BansalDoesTheWholeExceedItsParts; 2021ToTrustOrToThink), we took steps to neutralize trust effects so that observed differences in reliance could be attributed primarily to the visualizations. Specifically, we avoided anthropomorphizing the AI (Kevin2015TrustInAuto), consistently referring to it as “the AI”. We also described AI outputs as “suggestions” to counteract perfect automation schemas (Mary2002PerceivedUtility) and emphasize that these suggestions could be incorrect. Furthermore, we do not provide decision feedback for each trial (Dietvorst2014AlgorithmAP), or any information about the AI’s accuracy (Yin2019EffectofAccOnTrust).
@@ 4.8. Hypothesis
According to dual-process theory, human decision-making can be categorized into two distinct cognitive systems: rapid, heuristic-driven Type 1 thinking, and slower, analytical Type 2 thinking (Chaiklin2012ThinkingFA). A widely accepted distinction between these systems is that Type 2 thinking demands significantly more cognitive resources than Type 1 (Chaiklin2012ThinkingFA).
Under conditions of high time pressure, users often lack sufficient cognitive bandwidth to analytically evaluate AI-generated suggestions, potentially resulting in inappropriate reliance, either rejecting correct suggestions or accepting suboptimal ones (Lfstrm2023OnTD). This risk is especially pronounced with 2D minimaps, as these interfaces require users to mentally map the information to their corresponding physical environment, imposing substantial cognitive load due to reference-frame switching (Harris2013UnderstandingST). In contrast, embedded visualizations can substantially reduce or eliminate these spatial mapping demands by directly fusing information within the physical environment, thus potentially fostering more appropriate reliance on AI suggestions.
@@ 5.1. Quantitative Results
Prior to analysis, we excluded trials in which participants failed to make a decision within 20 seconds. After filtering, we retained 248 trials in Minimap+AI, 247 in Minimap+NoAI, 237 in X-ray+AI, and 237 in X-ray+NoAI, out of 1024 (32 participants $\times$ 32 trials). For each measure, trials were aggregated by participant and conditions being analyzed. We first tested for normality using the Shapiro-Wilk test. Results indicated that accuracy, point error, and AI reliance significantly violated the assumption of normality, whereas time and confidence did not. Accordingly, we used repeated-measures ANOVA ($\alpha=.05$) to assess the significance effects on time and confidence; then, post-hoc pairwise comparisons were conducted using Holm-Bonferroni corrections. For accuracy and point error, we used Aligned Rank Transform (ART) ANOVA (Wobbrock2011ART) to test significance and use ART-C (Elkin2021ART_C) for post-hoc contrast testing. To compare AI reliance in the Minimap+AI and X-ray+AI conditions, we used Wilcoxon signed-rank tests.
@@ 5.1.1. Decision Accuracy
Further observation suggests that X-ray+AI under-performance may stem from inappropriate reliance on AI. Participants in the X-ray+NoAI condition achieved a median decision accuracy of $Mdn=0.66$, which increased to $Mdn=0.70$ in the X-ray+AI condition. However, this improvement still fell short of the AI standalone baseline accuracy of 0.75, suggesting that participants may have relied on it inappropriately. This aligns with findings in prior empirical studies (2021BansalDoesTheWholeExceedItsParts). In contrast, performance in both Minimap+NoAI ($Mdn=0.88$) and Minimap+AI ($Mdn=0.88$) conditions exceeded the AI baseline, indicating that participants were able to appropriately rely on AI by identifying and overriding poor suggestions.
@@ 5.1.3. Response Time
Response Time Excluding Initial Search. We observed that participants in the X-ray conditions spent additional time to visually search the spatially distributed targets—especially those outside their immediate field of view. In contrast, in the Minimap conditions, all targets were presented in a 2D overview, making search overhead negligible. Prior decision-making model (Simon’s model (Simon1977)) and empirical studies (reisen2008identifying; russo1994eye) suggest that actual decision-making begins only after sufficient information has been gathered. Thus, similar to prior research (Jakobsen2013VISonDisplay), we distinguished the initial search phase from the subsequent decision-making phase. This motivates us to exclude the initial search overhead in the X-ray conditions to better compare its decision-making phase with the Minimap conditions.
@@
Following prior visualization-decision work (Padilla2018DecisionMW), we consider that meaningful decision-making start only after at least two options have been identified. Thus, for each trial in the X-ray conditions, we excluded the initial search time required to discover the first two distinct targets. Specifically, we defined two search segments to exclude: S1-from the start of the trial to the first frame when the first target entered the user’s field of view; and S2-from the frame when the user turned away from the first target (Figure 9a) to the frame when the second target appeared (Figure 9b). If a target was already visible at trial start, or if the first and second targets appeared simultaneously, the corresponding segment (S1 or S2) was set to zero. This definition provides a strict lower bound on search time, as participants may still require additional head movement to center the target in view. Yet, we acknowledge that this measure is not a theoretically perfect indicator of true decision time. Nevertheless, it provides a scientifically reasonable approximation that is consistent with definitions and procedures used in prior research (Jakobsen2013VISonDisplay). We discuss its limitations and possible refinements in Sec. 6.5.
Two co-authors independently annotated these segments frame by frame from the first-person recordings to ensure inter-rater reliability. Inter-rater reliability was assessed using a two-way mixed effects intraclass correlation coefficient, ICC(3,1), which indicated excellent agreement ($\mathrm{ICC}(3,1)=0.995$, 95% CI [0.992, 0.997]) (koo2016guideline). Encoding consistency was ensured through frame-level accurate marking and the clear visual cues afforded by first-person video recordings. Discrepancies were discussed and resolved through consensus. After subtracting the annotated initial search time from the response time, we observed a different pattern (Figure 7c). The analysis showed that there is a significant effect of Visualizations $\times$ AI Conditions interaction ($F(1,31)=7.70,p=\text{.009}$). Further post-hoc analysis indicated that participants in X-ray+AI ($M=11.23,SE=0.48$) spent less time than those in Minimap+AI ($M=12.08,SE=0.48$) with a significant effect ($\Delta_{M}=\text{-0.85},F(1,31)=-3.27,p=\text{.016}$), indicating that, with AI assistance, decisions were made more quickly with the X-ray visualization. No significant difference was observed between Minimap+NoAI and X-ray+NoAI, suggesting that, without AI assistance, participants spent comparable time making decisions across visualizations. In summary, although raw response times were longer in X-ray conditions, a detailed breakdown suggests that this extra time wasn’t entirely spent on the decision-making. This leads to interesting implications that we will further discuss in Sec. 6.4.

### v4

@@ 3.1. Task Abstraction and Definition
Spatial decision-making has long been studied under the broader umbrella of wayfinding in spatial cognition (Emo et al., 2012; Wiener et al., 2009). Foundational work by Siegel and White (Siegel and White, 1975) conceptualizes wayfinding as a decision process involving three forms of spatial knowledge: landmark, route, and survey. Motivated by real-world applications, our study particularly focuses on target selection (landmark). In tasks such as exploring a building, security surveillance, and emergency response, spatial decision-making often involves identifying or selecting a target before proceeding with an action like navigation or intervention. Although AR systems have frequently been investigated for supporting the action phase such as route guidance (Xu et al., 2024a), their potential to enhance the preceding decision-making step remains underexplored.
@@
  • Space - Large-scale Indoor Environment: These tasks often take place in physical spaces larger than the immediate space around a person such as buildings or open areas (Giuliano et al., 2013; Guerrieri et al., 2006). Tversky (Tversky, 2003) characterize such spaces as space of navigation. In this study, we focus on indoor building environments, given their prevalence across numerous application domains.
  • Time - Sensitive: Many high-stakes spatial tasks impose stringent time constraints, requiring users to make optimal decision as soon as possible (Nydegger et al., 2011). In this study, we focus on time-pressure scenarios that particularly benefit from decision support systems.
  • Data - Static and Dynamic: Appropriate target selection relies on both static data like building level layouts, and dynamic updates from sensors like walking people, occupant counts, and hazard alerts to evaluate each target (Adib and Katabi, 2013).
  • Decision - Multi-Criteria: These tasks frequently require decisions among multiple targets or locations distributed over rooms, floors, or areas (Launder and Perry, 2014). Success often involves balancing or prioritizing several factors, such as response time, travel distance, and urgency (Penney et al., 2022).
@@
X-Ray. Given our task focuses on large-scale indoor environments, we select AR X-ray (Figure 1 Left) as our embedded visualizations, which overlays a 1:1 scale 3D map on the real-world. It allows users see the interior layouts or objects behind walls, as if they had “X-ray” vision (Livingston et al., 2013), which has been widely used in applications like navigation (Xu et al., 2024b) and construction (Soria et al., 2018). For our study, we follow previous similar research (Xu et al., 2024b) and portray the building’s indoor layout, along with multiple targets (represented as cubes), and their associated data (the number of people in line) using virtual avatars.
Alternatives. World-in-Miniature (WiM) (Stoakley et al., 1995; Danyluk et al., 2021) provides a three-dimensional representation of the environment and can be implemented on flat-screen or AR displays. However, WiM typically relies heavily on interactions such as panning, tilting, zooming, or rotation to fully leverage their advantages. Because our current investigation prioritizes evaluating the impact of the visual representations themselves (without the influence of interactive complexity), we intentionally avoided such interactions in our study. Thus, we reserve exploration of interactive WiM and other advanced visualization methods for future research.
@@ 3.3. Visual Decision Supports
A recent survey by Oral et al. (Oral et al., 2023) highlights a broad spectrum of visual decision support tools, ranging from basic information visualizations to advanced AI-based suggestions (Lawless et al., 2024). For this study, we specifically choose to implement two fundamental yet representative forms of decision support across both the Minimap and X-ray:
@@
  • AI-Suggestions that highlight a single optimal target based on predefined criteria, with no accompanying explanation. Similar to all AI systems, suggestions may not always be accurate. We detail our adjustment of the performance of the AI based on prior works (Swaroop et al., 2024) and pilot studies in Sec. 4.7.
Following practices established in prior research (Vasconcelos et al., 2023), we deliberately constrain our decision support to these fundamental features for several reasons: First, target and real-time data visualization represent core functionalities broadly used in existing decision-support tools. Second, simple AI suggestions enable direct investigation of user reliance on AI assistance under different visualizations, without the additional cognitive complexity and potential confounding effects introduced by detailed explanations (Jiang et al., 2018).
While more advanced support, such as comparative visualizations, AI-generated explanations, and uncertainty displays, may offer added benefits, we intentionally focus on these fundamental forms to isolate the role of embedded visualizations in spatial decision-making and AI reliance. This controlled setup provides a baseline for future studies exploring richer decision supports.
@@ 4.2. Spatial Arrangement
An example spatial arrangement of the coffee machine targets locations for a single participant location, illustrating four distinct difficulty levels: same+close, same+far, cross+close, and cross+far. For each participant location, there are 4 trials, one difficulty level per trial. The participant’s location is marked in red, and colored dots indicate target locations in each difficulty level. The white dashed lines indicate that the coffee machine targets are placed at equal line-of-sight distances from the participant.
@@
The 8 participant locations in the study, distributed across two floors of the building. Participants were instructed to complete 4 trials at each location, following the order indicated by the numbered labels and route. The letter “L” is an abbreviation for location.
@@
Since the values of $d$ were fixed by the spatial arrangement of the targets, we manipulated the queue lengths $n$ to differentiate target optimality. Following prior research (Swaroop et al., 2024), each trial was designed to include intentionally varied options: one optimal, two suboptimal, and one worst choice, differing by at least 10 seconds in $T_{\text{coffee}}$. This setup allowed us to evaluate potential participant over-reliance on AI suggestions. Detailed calculations are provided in the supplementary materials.
@@ 4.3. Implementation and Apparatus
Following similar work (Buçinca et al., 2021), a simulated AI was used to provide decision suggestions in applicable trials. Details of the simulated AI setup are provided in Sec. 4.7. We developed the minimap based on high-precision CAD models provided by the building manager. The AR visualizations was developed using ARKit  [^3] https://developer.apple.com/augmented-reality/arkit/ and AR Mesh Manager   [^4] https://docs.unity3d.com/Packages/com.unity.xr.arkit@6.2/manual/arkit-meshing.html to scan the building and construct a 3D digital twin in Unity 6000.0.29f1, with ARAnchor Manager   [^5] https://docs.unity3d.com/Packages/com.unity.xr.arfoundation@6.2/manual/features/anchors/aranchormanager.html ensuring proper registration of the digital model with the physical environment. The system is available at our public repository https://github.com/demoPlz/ARAIReliance. All trials were completed using an Apple Vision Pro to eliminate device-related differences and provided a high-fidelity immersive experience.
@@ 4.6. Measures
For every trial, we documented and calculated the following measures based on prior research (Swaroop et al., 2024; Salimzadeh et al., 2024):
  • Decision Accuracy: We defined decision accuracy as the average point score across trials in each condition, ranging from 0 to 1. Following previous research (Swaroop et al., 2024), participants earned points according to their choice quality: optimal choice (1 point), suboptimal choices (0.5 points), and the worst choice (0 points). Each trial contained exactly one optimal target, two suboptimal targets, and one worst target.
  • AI Reliance: Following previous studies (Yang et al., 2020), we categorized participant reliance on AI as either appropriate or inappropriate (Figure 6). We classified reliance as appropriate when participants either followed correct AI suggestions or overrode incorrect ones with better choices. Inappropriate reliance included: (1) over-reliance: accepting suboptimal or worst suggestions; and (2) under-reliance: rejecting correct suggestions in favor of a worse option.
@@
  • Point Error: Similar to prior research (Hu et al., 2021; Warden et al., 2022), we assessed whether participants correctly pointed to the spatial location of their selected target based on first-person video recordings. Two co-authors independently coded each response as 0 (correct) or 1 (incorrect) to ensure inter-rater reliability. Inter-rater reliability was evaluated via Cohen’s Kappa ($\kappa=0.898$), suggesting almost perfect agreement according to Landis and Koch’s interpretation (Landis and Koch, 1977). Discrepancies were discussed and resolved through consensus. This measure indicates if participants actually knew the spatial location of their selected target.
@@ 4.7. Using Simulated AI
Similar to prior works (Swaroop et al., 2024; Buçinca et al., 2021), we used a simulated AI with an average accuracy of 0.75 to ensure controlled experimentation. In each AI condition, the AI suggested the optimal target in five trials, a suboptimal target in two trials, and the worst target in one trial (Figure 5). This setup yields an average accuracy of 0.75 = $\frac{5\times 1+2\times 0.5+1\times 0}{8}$. The order of these suggestions was randomized and counterbalanced across participants, ensuring that optimal, suboptimal, and worst suggestions were evenly distributed across all trials and participants.
Mitigating Effects of Trust in AI. Following previous work (Bansal et al., 2021; Buçinca et al., 2021), we took steps to neutralize trust effects so that observed differences in reliance could be attributed primarily to the visualizations. Specifically, we avoided anthropomorphizing the AI (Hoff and Bashir, 2015), consistently referring to it as “the AI”. We also described AI outputs as “suggestions” to counteract perfect automation schemas (Dzindolet et al., 2002) and emphasize that these suggestions could be incorrect. Furthermore, we do not provide decision feedback for each trial (Dietvorst et al., 2014), or any information about the AI’s accuracy (Yin et al., 2019).
@@ 4.8. Hypothesis
According to dual-process theory, human decision-making can be categorized into two distinct cognitive systems: rapid, heuristic-driven Type 1 thinking, and slower, analytical Type 2 thinking (Chaiklin, 2012). A widely accepted distinction between these systems is that Type 2 thinking demands significantly more cognitive resources than Type 1 (Chaiklin, 2012).
Under conditions of high time pressure, users often lack sufficient cognitive bandwidth to analytically evaluate AI-generated suggestions, potentially resulting in inappropriate reliance, either rejecting correct suggestions or accepting suboptimal ones (Löfström, 2023). This risk is especially pronounced with 2D minimaps, as these interfaces require users to mentally map the information to their corresponding physical environment, imposing substantial cognitive load due to reference-frame switching (Harris et al., 2013). In contrast, embedded visualizations can substantially reduce or eliminate these spatial mapping demands by directly fusing information within the physical environment, thus potentially fostering more appropriate reliance on AI suggestions.
@@ 5.1. Quantitative Results
Prior to analysis, we excluded trials in which participants failed to make a decision within 20 seconds. After filtering, we retained 248 trials in Minimap+AI, 247 in Minimap+NoAI, 237 in X-ray+AI, and 237 in X-ray+NoAI, out of 1024 (32 participants $\times$ 32 trials). For each measure, trials were aggregated by participant and conditions being analyzed. We first tested for normality using the Shapiro-Wilk test. Results indicated that accuracy, point error, and AI reliance significantly violated the assumption of normality, whereas time and confidence did not. Accordingly, we used repeated-measures ANOVA ($\alpha=.05$) to assess the significance effects on time and confidence; then, post-hoc pairwise comparisons were conducted using Holm-Bonferroni corrections. For accuracy and point error, we used Aligned Rank Transform (ART) ANOVA (Wobbrock et al., 2011) to test significance and use ART-C (Elkin et al., 2021) for post-hoc contrast testing. To compare AI reliance in the Minimap+AI and X-ray+AI conditions, we used Wilcoxon signed-rank tests.
@@ 5.1.1. Decision Accuracy
Further observation suggests that X-ray+AI under-performance may stem from inappropriate reliance on AI. Participants in the X-ray+NoAI condition achieved a median decision accuracy of $Mdn=0.66$, which increased to $Mdn=0.70$ in the X-ray+AI condition. However, this improvement still fell short of the AI standalone baseline accuracy of 0.75, suggesting that participants may have relied on it inappropriately. This aligns with findings in prior empirical studies (Bansal et al., 2021). In contrast, performance in both Minimap+NoAI ($Mdn=0.88$) and Minimap+AI ($Mdn=0.88$) conditions exceeded the AI baseline, indicating that participants were able to appropriately rely on AI by identifying and overriding poor suggestions.
@@ 5.1.3. Response Time
Response Time Excluding Initial Search. We observed that participants in the X-ray conditions spent additional time to visually search the spatially distributed targets—especially those outside their immediate field of view. In contrast, in the Minimap conditions, all targets were presented in a 2D overview, making search overhead negligible. Prior decision-making model (Simon’s model (Simon, 1977)) and empirical studies  (Reisen et al., 2008; Russo and Leclerc, 1994) suggest that actual decision-making begins only after sufficient information has been gathered. Thus, similar to prior research  (Jakobsen and Hornbæk, 2013), we distinguished the initial search phase from the subsequent decision-making phase. This motivates us to exclude the initial search overhead in the X-ray conditions to better compare its decision-making phase with the Minimap conditions.
@@
Following prior visualization-decision work (Padilla et al., 2018b), we consider that meaningful decision-making start only after at least two options have been identified. Thus, for each trial in the X-ray conditions, we excluded the initial search time required to discover the first two distinct targets. Specifically, we defined two search segments to exclude: S1-from the start of the trial to the first frame when the first target entered the user’s field of view; and S2-from the frame when the user turned away from the first target (Figure 9a) to the frame when the second target appeared (Figure 9b). If a target was already visible at trial start, or if the first and second targets appeared simultaneously, the corresponding segment (S1 or S2) was set to zero. This definition provides a strict lower bound on search time, as participants may still require additional head movement to center the target in view. Yet, we acknowledge that this measure is not a theoretically perfect indicator of true decision time. Nevertheless, it provides a scientifically reasonable approximation that is consistent with prior research  (Jakobsen and Hornbæk, 2013). We discuss its limitations and possible refinements in Sec. 6.5.
Two co-authors independently annotated these segments frame by frame from the first-person recordings to ensure inter-rater reliability. Inter-rater reliability was assessed using a two-way mixed effects intraclass correlation coefficient, ICC(3,1), which indicated excellent agreement ($\mathrm{ICC}(3,1)=0.995$, 95% CI [0.992, 0.997]) (Koo and Li, 2016). Encoding consistency was ensured through frame-level accurate marking and the clear visual cues afforded by first-person video recordings. Discrepancies were discussed and resolved through consensus. After subtracting the annotated initial search time from the response time, we observed a different pattern (Figure 7c). The analysis showed that there is a significant effect of Visualizations $\times$ AI Conditions interaction ($F(1,31)=7.70,p=\text{.009}$). Further post-hoc analysis indicated that participants in X-ray+AI ($M=11.23,SE=0.48$) spent less time than those in Minimap+AI ($M=12.08,SE=0.48$) with a significant effect ($\Delta_{M}=\text{-0.85},F(1,31)=-3.27,p=\text{.016}$), indicating that, with AI assistance, decisions were made more quickly with the X-ray visualization. No significant difference was observed between Minimap+NoAI and X-ray+NoAI, suggesting that, without AI assistance, participants spent comparable time making decisions across visualizations. In summary, although raw response times were longer in X-ray conditions, a detailed breakdown suggests that this extra time wasn’t entirely spent on the decision-making. This leads to interesting implications that we will further discuss in Sec. 6.4.